OpenStudy (anonymous):

Use the square root property to solve the equation: 3x^2-39=0

OpenStudy (anonymous):

3x^2 = 39 x^2 = 13 x = SQRT(13)

OpenStudy (anonymous):

\[\LARGE 3x^2-39=0\] \[\LARGE 3x^2=39\] \[\LARGE x^2=\frac{39}{3}\] \[\LARGE x^2=3\] \[\LARGE x=\pm\sqrt{3}\]

OpenStudy (anonymous):

how would i type that as a solution set?

OpenStudy (anonymous):

\[3x ^{2}-39=0\] \[3x ^{2}=39\] \[x ^{2}=13\] x=\[\sqrt{13} or x= \sqrt[-]{13}\] now what do i do?

OpenStudy (anonymous):

use \pm for \[\LARGE \pm\]

OpenStudy (anonymous):

no im saying how do i type that as an ordered pair, is it \[(\sqrt{13},-\sqrt{13}) ??\]

OpenStudy (anonymous):

\[\LARGE x\in \{\sqrt{3},-\sqrt{3}\}\] \LARGE x\in \{ \sqrt{3} , -\sqrt{3} \} this is how you type in Equation button !

OpenStudy (anonymous):

sorry... I made a mistake in my TWO postst replace \sqrt{13} instead of \sqrt{3}

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you're welcome :)